Skip to content
CalcGospel 國際數學圖譜
返回

CIE 9709 2026 June Paper 55 Q3

A Level / CIE / S1

CIE 9709 2026 June Paper 55 Paper · Question 3

题目

Problem

For her morning break, Jules has a drink. She has tea, juice or water, with probabilities 0.2, 0.35 and 0.45 respectively.

When she has tea, the probability that she has a cake is 0.25.

When she has juice, the probability that she has a cake is 0.4.

When she has water, the probability that she has a cake is xx.

On a randomly chosen day, the probability that Jules has a cake is 0.505.

(a) Find the value of xx.

(2)

(b) Find the probability that Jules has juice given that she does not have a cake.

(3)
题目中文翻译

Jules 在上午休息时会喝一种饮品。她喝茶、果汁或水的概率分别为 0.2、0.35 和 0.45。

当她喝茶时,她吃蛋糕的概率为 0.25。

当她喝果汁时,她吃蛋糕的概率为 0.4。

当她喝水时,她吃蛋糕的概率为 xx。

随机选取一天,Jules 吃蛋糕的概率为 0.505。

(a) 求 xx 的值。

(2)

(b) 已知 Jules 没有吃蛋糕,求她喝果汁的概率。

(3)
原卷通用作答要求

Answer all questions. You should use a calculator where appropriate.

You must show all necessary working clearly; no marks will be given for unsupported answers from a calculator.

Give non-exact numerical answers correct to 3 significant figures, or 1 decimal place for angles in degrees, unless a different level of accuracy is specified in the question.

请回答所有题目。适当时应使用计算器。必须清楚展示所有必要的计算过程;仅给出计算器答案而没有过程,不予评分。除非题目另有精度要求,非精确数值答案保留 3 位有效数字,以度为单位的角度保留 1 位小数。

解答

(a)

解法一

思路

展开

分别计算饮茶、果汁和水且吃蛋糕的概率,三个互斥分支相加等于吃蛋糕的总概率。

答题过程

展开

Let CC denote having a cake. By conditioning on the drink,

P(C)= 0.2(0.25)+0.35(0.4)+0.45x0.505= 0.05+0.14+0.45x,0.45x= 0.315,x= 0.7.\begin{align*} \mathrm{P}(C)=&\,0.2(0.25)+0.35(0.4)+0.45x\\[4mm] 0.505=&\,0.05+0.14+0.45x,\\[4mm] 0.45x=&\,0.315,\\[4mm] x=&\,\boxed{0.7}. \end{align*}

(b)

解法一

思路

展开

分子是喝果汁且不吃蛋糕,分母是所有不吃蛋糕的概率。果汁分支中不吃蛋糕的条件概率为0.6。

答题过程

展开

Let JJ denote having juice. Then

P(J∩C′)=0.35(1−0.4)=0.21,\mathrm{P}(J\cap C')=0.35(1-0.4)=0.21,

and P(C′)=1−0.505=0.495\mathrm{P}(C')=1-0.505=0.495. Therefore,

P(J∣C′)=0.210.495=1433.\mathrm{P}(J\mid C')=\frac{0.21}{0.495} =\boxed{\frac{14}{33}}.